Im and the region of Ro RG 15V 1-(50) Find: Ives Vas Im geration. km = 60 mAlv? y = 2V Ra IMR RI 200.12 ir and the region
Answer #1

Solution 15V mo RO RG VGx ID D IG=0 G. S -> 16=0 = Voltage drop across Rg is tero C VG Foo VG = VD E Vp > VG - Vy → Vos 2 VGSVG = VD Io = kn CVG - KM ID = kn CVD-VA) (2) from o § 6 15-V) = Kn CVD-V) 200 15- Vo = (60m) (200) ( VD V3-2 VD VY) 15-vo =